OpenStudy (anonymous):

This is an Geometric Problem A lot is in the shape of a triangle . One side is 32 feet longer than the shortest side while the third side is ten feet shorter than twice the shorter side . The perimeter of the lot is 238 feet . Find the length of the sides of the lot .

OpenStudy (anonymous):

i think it is just an algebraic problem :) lolz let the shortest side be x then try to write all other sides in terms of x then sum of all is given 238 feet so x can be solved :)

OpenStudy (anonymous):

try if any problem persist i will help you :)

OpenStudy (anonymous):

can you solve this question ?

OpenStudy (anonymous):

smallest side=x one side=x+32 other side=2x-10 so sum of all side i.e. perimeter=x+x+32+2x-10=4x+22 but sum is given=238 4x+22=238 4x=238-22=216 x=216/4=54feet so smallest side=54 feet one side=54+32=86 feet other=2*54-10=108-10=98 feet :)

OpenStudy (anonymous):

so the answer is 98 feet ...

OpenStudy (dwelsher):

x=shortest side x+32=another side (2x-10)=final side perimeter = 238 so--x+x+32+2x-10=238 4x=213 x=54,,,,x+32=86,,,,2x-10=98...........54+86+98=238